Al. Let T1(x, y, z) = (1-y+z, 2:0 – y + 2z, 2y + 2). (a). Is T1 one-to-one?

(b). Is T onto?
Answer #1

A1. T1: R3 →R3 is defined by T1(x,y,z) = (x-y+z, 2x-y+2z, 2y+z).

Hence T1 (e1) = T1 (1,0,0) = (1,2,0), T1 (e2) = T1 (0,1,0) = (-1,-1,2) and T1 (e3) = T1 (0,0,1) = (1,2,1).

Hence the standard matrix of T1 is A(say) = [T1 (e1), T1 (e2), T1 (e3)] =

1

-1

1

2

-1

2

0

2

1

The RREF of A is

1

0

0

0

1

0

0

0

1

This implies that the columns of A are linearly independent and span R3.

Therefore,

(a). T1 is one-to-one.

(b) T1 is onto.